OpenStudy (jojo12g):

What is the sign of the answer? (1.3) • (–4) • (–0.8) • (14) • (–6) A. positive B. negative

OpenStudy (just_one_last_goodbye):

You tell me http://i.imgur.com/u03zTZ2.png

OpenStudy (anonymous):

positive

OpenStudy (just_one_last_goodbye):

^ LOL XD wow...

OpenStudy (just_one_last_goodbye):

@jojo12g if you get this one wrong it's cause you want to fail

OpenStudy (calculusxy):

let's review some basics \(+ \times + = +\) \(+ \times - = -\) \(- \times - = +\) \(- \times + = -\) break this down into negatives and positives together. (1.3) • (–4) • (–0.8) • (14) • (–6) -> Negatives: \(-4 \times -0.8\times -6\) which means a negative because if you have an odd number of negatives, it means a negative. Positives: \(1.3 \times 14\) and that is a positive because \(+ \times + = +\) now we have a negative, \(- \), multiplied by a positive \(+\) and that would be \(- \times + = -\).

OpenStudy (anonymous):

The answer is positive since you're doing a sequence. A positive times a negative is a negative. Negative times a negative is a positive, positive times a positive is a positive, and positive times a negative is a negative. So the answer is B.) Negative.

OpenStudy (anonymous):

@just_one_last_goodbye why'd u say lol XD wow

OpenStudy (just_one_last_goodbye):

Although your breaking the CoC by giving direct answers, your answer is wrong ._. right after my steps showed the correct answer. the user can take the time to identify symbols's between negative and positive and tell the difference.